In Maryland, only three goods are produced domestically - popcorns, movie tickets, and diet drinks. The following table shows the price and quantities of these goods produced in the year 2018, 2019, 2020. 2018 is the base year



The market bundle for a typical family is 3 popcorns, 3 movie tickets and 5 diet drinks.



Calculate Nominal and Real GDP in all three years.


Calculate real economic growth rate in 2020


Calculate CPI in all three years.


Calculate inflation rate in 2020




2019



2020



2021





P



Q



P



Q



P



Q



Popcorn



1



500



1



600



1.05



590



Movie ticket



5



300



10



200



10.5



210



Diet drink



0.7



300



0.8



400



0.75



420

1
Expert's answer
2021-11-07T19:43:22-0500

Solution:

Popcorn:

1.). Nominal GDP 2019 = 500 x 1 = 500

Nominal GDP 2020 = 600 x 1 = 600

Nominal GDP 2021 = 590 x 1.05 = 619.5

Movie Tickets:

Nominal GDP 2019 = 300 x 5 = 1,500

Nominal GDP 2020 = 200 x 10 = 2,000

Nominal GDP 2021 = 210 x 10.5 = 2,205

Diet Drink:

Nominal GDP 2019 = 300 x 0.7 = 210

Nominal GDP 2020 = 400 x 0.8 = 320

Nominal GDP 2021 = 420 x 0.75 = 315

 

Popcorn:

Real GDP 2019 = 500 x 1 = 500

Real GDP 2020 = 600 x 1 = 600

Nominal GDP 2021 = 590 x 1 = 590

Movie Tickets:

Real GDP 2019 = 300 x 5 = 1,500

Real GDP 2020 = 200 x 5 = 1,000

Real GDP 2021 = 210 x 5 = 1,050

Diet Drink:

Real GDP 2019 = 300 x 0.7 = 210

Real GDP 2020 = 400 x 0.7 = 280

Real GDP 2021 = 420 x 0.7 = 294

 

2.). Real economic growth rate in 2020:

Popcorn = "(\frac{GDP\; in\; year\; 2020}{GDP\; in\; year\; 2019}) -1 = (\frac{600}{500}) - 1 = 20\%"


Movie ticket = "(\frac{GDP\; in\; year\; 2020}{GDP\; in\; year\; 2019}) -1 = (\frac{1000}{1500}) - 1 = 10 0\%"


Diet Drink = "(\frac{GDP\; in\; year\; 2020}{GDP\; in\; year\; 2019}) -1 = (\frac{280}{210}) - 1 = 33.33\%"

 

3.). CPI:

Cost of the market basket in base year (2019) = (1 x 500) + (5 x 300) + (0.7 x 300) = 2,210

Cost of the market basket in given year (2019) = (1 x 500) + (5 x 300) + (0.7 x 300) = 2,210

CPI 2019 = "\frac{2210}{2210} \times100 = 100\%"

Cost of the market basket in given year (2020) = (1 x 600) + (10 x 200) + (0.8 x 400) = 2,920

CPI 2020 = "\frac{2920}{2210} \times100 = 132.13\%" %

Cost of the market basket in given year (2021) = (1.05 x 590) + (10.5 x 210) + (0.75 x 420) = 3,139.5

CPI 2021 = "\frac{3139.5}{2210} \times100 = 142.06\%"

 

Inflation rate in 2020 = 132.13% - 100% = 32.13%